A rectangular steel sheet is bounded by the axis x = 0, y = 0, x = a and y = b. The temperature along the edge x = 0 are kept at 100°C and other edges are at 0°C. Let u(x,y) denote the temperature satisfying the equation [15] + d^2u/dx^2 + d^2u/dy^2 = 0 Find the steady state temperature u(x, y), by assuming the solution to be of the form u(x, y) = (AePX + Be-px)(C cospy + D sin py).


1
Expert's answer
2021-05-24T13:50:43-0400

Solution

Solution u(x, y) = (Aepx + Be-px)(C cospy + D sin py) satisfy the equation ∂2u/∂x2 + ∂2u/∂y2 = 0 because ∂2u/∂x2 = p2u, ∂2u/∂y2 = -p2u.

Satisfying boundary conditions at y = 0 and y = b we’ll get C = 0, D sin pb = 0. From the last equality p = πn/b for any natural n and the solution is

"u(x,y)\ =\ \sum_{n=1}^{\infty}{(A_ne^{\pi nx/b}\ +\ B_ne^{-\pi nx/b})}sin\left(\frac{\pi n}{b}y\right)"

Expanding function f(y) on interval (0,b) into the series by sin(πny/b)  we’ll get fn = 2[1-(-1)n]/ πn

Satisfying boundary conditions at x = 0 and x = a we’ll get

An + Bn = 200[1-(-1)n]/ πn, Aneπna/b + Bne-πna/b= 0   =>   Bn =  - Ane2πna/b , An (1- e2πna/b ) = 200[1-(-1)n]/ πn =>   An = 200[1-(-1)n]/ [πn (1- e2πna/b )], Bn =  - 200[1-(-1)n]e2πna/b / [πn (1- e2πna/b )].

Therefore solution is

"u(x,y)\ =\ \frac{200}{\pi}\sum_{n=1}^{\infty}\frac{e^{\pi nx/b}\ -\ e^{2\pi na/b}e^{-\pi nx/b}}{1-e^{2\pi na/b}}\frac{1-{(-1)}^n}{n}sin\left(\frac{\pi n}{b}y\right)\ =\"

"=\frac{400}{\pi}\sum_{m=0}^{\infty}\frac{sinh(\pi(2m+1)(a-x)/b)}{sinh(\pi(2m+1)a/b)}\frac{1}{(2m+1)}sin\left(\frac{\pi(2m+1)}{b}y\right)"

Answer

"u(x,y)=\frac{400}{\pi}\sum_{m=0}^{\infty}\frac{sinh(\pi(2m+1)(a-x)/b)}{sinh(\pi(2m+1)a/b)}\frac{1}{(2m+1)}sin\left(\frac{\pi(2m+1)}{b}y\right)"